OpenStudy (anonymous):

Solve: y²+400 = 0 A. y = -20i² B. y = ± 20 C. y = 20i D. y = ± 20i

OpenStudy (anonymous):

B is correct

OpenStudy (anonymous):

Thanks. Can you help me with a similar one?

OpenStudy (anonymous):

This one: Solve: 4z² + 39 = 7 A. -4 B. 4i C. ±2isqrt2 D. 2isqrt2 sqrt= square root

OpenStudy (anonymous):

4z^2 = -32 z^2= -8 z=+- 2i sqrt 2 so B is correct